### v1

# Language Model with Retrieval-Augmented Generation

Abstract
Embedding-based retrieval models have made significant strides in retrieval-augmented generation (RAG) techniques for text and multimodal large language models (LLMs) applications. However, when it comes to speech larage language models (SLLMs), these methods are limited to a two-stage process, where automatic speech recognition (ASR) is combined with text-based retrieval. This sequential architecture suffers from high latency and error propagation. To address these limitations, we propose a unified embedding framework that eliminates the need for intermediate text representations. Specifically, the framework includes separate speech and text encoders, followed by a shared scaling layer that maps both modalities into a common embedding space. Our model reduces pipeline latency by 50% while achieving higher retrieval accuracy compared to traditional two-stage methods. We also provide a theoretical analysis of the challenges inherent in end-to-end speech retrieval and introduce architectural principles for effective speech-to-document matching. Extensive experiments demonstrate the robustness of our approach across diverse acoustic conditions and speaker variations, paving the way for a new paradigm in multimodal SLLMs retrieval systems.
1 Introduction
In recent years, embedding models have made remarkable strides, becoming foundational components for a wide range of natural language processing tasks(Su et al., 2023) (Xiao et al., 2024) (Wang et al., 2023). These models excel at capturing semantic relationships within text data, enabling robust retrieval and matching capabilities. The success of text embeddings has naturally extended to explorations in multimodal domains, leading to significant advancements in text-image embedding models that effectively align semantic spaces across modalities (Chen et al., 2022) (Yasunaga et al., 2023) (Hu et al., 2023). However, despite the pervasive presence of speech content in real-world applications, the integration of speech modalities into embedding-based retrieval systems remains underexplored, highlighting a critical gap in multimodal information processing.
Figure 1: The traditional two-stage speech retrieval system and our proposed unified speech-text embedding based retrieval system, where * indicates that some modules in the model can be shared.
This challenge becomes particularly pronounced in speech-based retrieval-augmented generation (RAG) systems, which have emerged as a promising approach to augment speech large language models (SLLMs) with external knowledge (Jacqmin et al., 2023). Current speech RAG systems predominantly adopt a two-stage paradigm: automatic speech recognition (ASR) followed by text-based RAG as shown in Fig 1. While this approach leverages well-established components, it suffers from two major limitations. First, the sequential processing of ASR and text-based RAG significantly increases system latency, rendering it impractical for real-time applications where rapid responses are critical. Second, and more fundamentally, this cascaded architecture is prone to error propagation. ASR errors—particularly in challenging acoustic environments or with non-standard speech—inevitably degrade downstream retrieval performance. This cascading effect can result in entirely irrelevant document retrieval, even when ASR errors impact only a few key terms.
Several attempts have been made to address these challenges through intermediate solutions, such as optimizing ASR for retrieval-specific metrics or implementing early fusion techniques (Serdyuk et al., 2018) (Szymański et al., 2023). However, these approaches still maintain the fundamental separation between speech processing and document retrieval, failing to capture the direct relationship between acoustic patterns and document semantics. Furthermore, existing methods often struggle with the inherent temporal nature of speech and the challenge of aligning variable-length speech sequences with document embeddings.
To overcome these limitations, we propose an end-to-end speech RAG model that directly learns to map speech inputs to relevant document embeddings, enhancing the performance and scalability of SLLMs. Our approach begins with separate speech and text encoders to extract features from each modality, followed by a shared scaling layer to produce the final unified speech-text embeddings. By learning a unified embedding space that captures both acoustic and semantic features, our method eliminates the need for intermediate text representations. Through careful architectural design and innovative training strategies, we achieve robust speech-to-document matching which is resilient to acoustic variations and speaker differences. Extensive experiments demonstrate that our model not only reduces pipeline latency by 50% but also achieves higher retrieval accuracy compared to traditional two-stage systems. Our contributions open new possibilities for real-time speech-based information retrieval and challenge the conventional wisdom about the necessity of intermediate text representations in processing pipelines of RAG based SLLMs. The proposed approach not only advances the state-of-the-art in speech RAG systems but also provides a foundation for future research in end-to-end multimodal retrieval systems.
2 Related Works
2.1 Multimodal Large Language Models
Recent advances in multimodal large language models (MLLMs) (Li et al., 2022) (Li et al., 2023) (Liu et al., 2024) have demonstrated impressive capabilities in processing visual and linguistic information. Models like LLaVA (Liu et al., 2024) have established effective architectures combining LLMs with visual encoders through carefully designed projection layers. This architecture typically involves a two-stage training process: first aligning visual-textual representations, then fine-tuning on instruction data. While most efforts have focused on vision-language models, some recent work has explored speech-language models by adapting similar architectures (Zhang et al., 2023) (Fang et al., 2024). However, these models primarily focus on speech understanding and generation rather than creating robust speech embeddings for retrieval tasks. Our work bridges this gap by introducing an embedding-focused architecture specifically designed for speech-text alignment.
2.2 Multimodal Embeddings
In the field of multimodal embeddings, CLIP (Radford et al., 2021) has established a strong foundation for vision-language understanding through contrastive learning on large-scale image-text pairs. Its success has inspired numerous follow-up works in visual-textual alignment and retrieval. While vision-language embedding models have made significant progress, the speech domain remains relatively unexplored. Current speech-text retrieval systems typically rely on a cascaded approach, first converting speech to transcripts through ASR Hinton et al. (2012); Prabhavalkar et al. (2024); Zhang et al. (2024) before applying text-based retrieval methods. This two-stage process, while straightforward, introduces latency and error propagation issues. The lack of end-to-end speech embedding models, analogous to CLIP’s role in vision, represents a significant gap in multimodal understanding. Our work addresses this limitation by introducing a unified speech embedding framework that directly aligns speech with textual content, demonstrating that end-to-end speech-text embeddings can achieve superior performance while reducing computational overhead.
Figure 2: Overview of Our methods. The stage 1 is align training and the stage 2 is contrastive training
3 Method
In this section, we introduce our proposed method (Fig 2) and the two baseline methods we experimented with.
3.1 Training Workflow
3.1.1 Speech-Text Alignment Pre-Training
We propose a two-stage training approach, with pre-training focusing on aligning speech and text embeddings in a unified semantic space. Given an input pair of speech signal $x_{a}$ and its corresponding text transcript $x_{t}$, we first extract their respective features:
$h_{s}=f_{s}(x_{s})\in\mathbb{R}^{T\times d_{s}}$ (1)
$z_{t}=f_{t}(x_{t})\in\mathbb{R}^{L\times d_{t}}$ (2)
where $f_{s}$ and $f_{t}$ denote the speech and text encoders. $T$ and $L$ denote the sequence lengths of speech and text features respectively, while $d_{s}$ and $d_{t}$ represent their feature dimensions.
To bridge the modality gap between speech and text features, we introduce an adaptation module inspired by LLaVA’s (Liu et al., 2024) design. This module consists of two components: a temporal convolutional layer for dimension reduction and a Multi-Layer Perceptron (MLP) for feature projection. The adaptation process can be formulated as:
$z_{s}=\text{MLP}(\text{Conv1D}(h_{s}))\in\mathbb{R}^{T^{\prime}\times d_{t}}$ (3)
where $T^{\prime}$ is the reduced sequence length after convolution, and the output dimension matches BERT’s(Devlin, 2018) hidden size $d_{t}$. The Conv1D layer uses a kernel size of k and stride s to effectively capture local temporal dependencies while reducing the sequence length. The MLP consists of two linear layers with a GELU activation function:
$\text{MLP}(x)=W_{2}(\text{GELU}(W_{1}x+b_{1}))+b_{2}$ (4)
During pre-training, we minimize the Mean Squared Error (MSE) loss between the adapted speech features and text embeddings:
$\mathcal{L}_{\text{pre}}=\big{\|}\frac{1}{T^{\prime}}\sum_{i=1}^{T^{\prime}}z_{s}^{i}-\frac{1}{L}\sum_{i=1}^{L}z_{t}^{i}\big{\|}_{2}^{2}$ (5)
where $z_{s}^{i}$ and $z_{t}^{i}$ represent the i-th token embeddings of the adapted speech and text features respectively. This alignment objective ensures that the speech features capture semantic information that corresponds well with textual representations. Our empirical analysis shows that this pre-training strategy effectively reduces the modality gap while maintaining the rich acoustic information necessary for accurate retrieval.
Algorithm 1 Cross-modal Speech-Text Alignment and Retrieval
0:  Speech signal $x_{s}$, text $x_{t}$, positive document $d^{+}$, negative documents $\{d^{-}_{i}\}_{i=1}^{N}$, temperature $\tau$
1:  // Stage 1: Pre-training
2:  $h_{s}\leftarrow f_{s}(x_{s})$
3:  $z_{t}\leftarrow f_{t}(x_{t})$
4:  $h_{s}^{\prime}\leftarrow\text{Conv1D}(h_{s})$
5:  $z_{s}\leftarrow\text{MLP}(h_{s}^{\prime})$
6:  $\mathcal{L}_{\text{pre}}\leftarrow\big{\|}\frac{1}{T^{\prime}}\sum_{i=1}^{T^{\prime}}z_{s}^{i}-\frac{1}{L}\sum_{i=1}^{L}z_{t}^{i}\big{\|}_{2}^{2}$
7:  // Stage 2: Task-specific Fine-tuning
8:  $q\leftarrow\text{Linear}(\text{Adapter}(f_{s}(x_{s})))$
9:  $k^{+}\leftarrow\text{Linear}(f_{t}(d^{+}))$
10:  for $i\leftarrow 1$ to $N$ do
11:     $k^{-}_{i}\leftarrow\text{Linear}(f_{t}(d^{-}_{i}))$
12:  end for
13:  // Select loss function by task type
14:  if task = retrieval then
15:     $\mathcal{L}\leftarrow-\log\frac{\exp(s(q,k^{+})/\tau)}{\exp(s(q,k^{+})/\tau)+\sum_{i=1}^{N}\exp(s(q,k^{-}_{i})/\tau)}$
16:  else if task = sts then
17:     $\mathcal{L}\leftarrow\log(1+\sum_{s_{ij}>s_{mn}}e^{(s_{mn}-s_{ij})/\tau})$
18:  else if task = classification then
19:     $\mathcal{L}\leftarrow-\log\frac{\exp(s(q,y^{+})/\tau)}{\exp(s(q,y^{+})/\tau)+\sum_{j}\exp(s(q,y^{-}_{j})/\tau)}$
20:  end if
21:  Update parameters using $\nabla\mathcal{L}_{\text{pre}}$ and $\nabla\mathcal{L}$
3.1.2 Speech-Text Retrieval Fine-Tuning
In the fine-tuning stage, we employ multi-task hybrid loss to optimize the alignment between speech queries and relevant document embeddings. For an speech query $x_{s}$ and its corresponding document $d^{+}$, along with a set of negative documents $\{d^{-}_{i}\}_{i=1}^{N}$, we first encode them using our pre-trained encoders. Inspired by text-embedding-v3 (OpenAI, 2024) and Piccolo2 (Huang et al., 2024), we scale up the embedding vector dimension to enhance model capacity. A shared linear layer is then added to the final layers of both the speech and text encoders. The operation can be formulated as:
$q=g_{\text{speech}}(x_{s})=\text{Linear}(\text{Adapter}(f_{s}(x_{s})))$ (6)
$k^{+}=g_{\text{doc}}(d^{+})=\text{Linear}(f_{t}(d^{+}))$ (7)
$k^{-}_{i}=g_{\text{doc}}(d^{-}_{i})=\text{Linear}(f_{t}(d^{-}_{i}))$ (8)
For different downstream tasks, we adopt different loss functions to better optimize various objectives:
For retrieval and reranking tasks, we use InfoNCE loss (Gutmann and Hyvärinen, 2010) with cosine similarity:
$\mathcal{L}_{\text{re}}=-\log\frac{\exp(s^{+}/\tau)}{\exp(s^{+}/\tau)+Z_{\text{neg}}}$ (9)
where $s^{+}=s(q,k^{+})$ denotes the cosine similarity between query and positive document, $Z_{\text{neg}}=\sum_{i=1}^{N}\exp(s(q,k^{-}_{i})/\tau)$ is the sum over negative pairs, and $\tau$ is the temperature parameter.
For STS and pair-classification tasks, considering the fine-grained nature of similarity labels, we employ the cosent loss:
$\mathcal{L}_{\text{sts}}=\log\left(1+\sum_{s_{ij}>s_{mn}}e^{z_{mn}/\tau}\right)$ (10)
where $s_{ij}=s(x_{i},x_{j})$ represents similarity score, $z_{mn}=\cos(x_{m},x_{n})-\cos(x_{i},x_{j})$.
For classification and clustering tasks, we reformulate the data into contrastive triplets using the SFR embedding method. Each input $x$ is paired with its target label $y^{+}$ as the positive pair, while the remaining labels $\{y^{-}_{j}\}$ serve as negative pairs:
$\mathcal{L}_{\text{cls}}=-\log\frac{\exp(\frac{s(x,y^{+})}{\tau})}{\exp(\frac{s(x,y^{+})}{\tau})+\sum_{j}\exp(\frac{s(x,y^{-}_{j})}{\tau})}$ (11)
The final loss function is dynamically selected based on the task type:
=$\mathcal{L}=\begin{cases}\mathcal{L}_{\text{cls}}&\text{if task is classification or clustering}\\ \mathcal{L}_{\text{sts}}&\text{if task is STS or pair-classification}\\ \mathcal{L}_{\text{re}}&\text{if task is retrieval or reranking}\end{cases}$ (12)
This multi-task hybrid loss approach enables our model to learn robust cross-modal representations that effectively capture the semantic relationships between speech queries and relevant documents across various downstream tasks.
3.2 Other baselines
3.2.1 Projection to text space
Following the design philosophy of vision-language models like LLaVA, we implement a baseline that projects speech features into text space through a carefully designed three-stage training process:
  • Stage 1: Train only the adapter module while freezing both speech and text encoders to establish initial cross-modal connections
  • Stage 2: Jointly optimize speech encoder and adapter while keeping text encoder fixed as stable semantic anchors
  • Stage 3: End-to-end training of the entire pipeline to refine cross-modal relationships
The speech features go through a speech-adapter-text encoding pipeline: $h=f_{t}(\text{Adapter}(f_{s}(x_{s})))$. While this approach successfully transfers vision-language strategies to speech domain, we find it challenging to preserve fine-grained acoustic information through the cascaded transformations, particularly for speech characteristics that lack direct textual correspondences.
3.2.2 Speech-Text Alignment with CTC Loss
We also explore using Connectionist Temporal Classification (CTC) loss (Graves et al., 2006) for temporal alignment between speech and text sequences:
$\mathcal{L}_{\text{CTC}}=-\log P(y|h_{\text{speech}})$ (13)
where $h_{\text{speech}}$ represents the adapted speech features and $y$ is the target text sequence. The adapter module maintains temporal information through strategically designed convolutional layers before projecting to the tokenizer’s index space of the text. While CTC provides explicit temporal supervision and theoretically enables better alignment between acoustic patterns and semantic content, our experiments reveal that its strong emphasis on frame-level alignment proves less effective for semantic retrieval tasks compared to our proposed approach.
Table 1: Performance Comparison of Different Methods on Various Tasks of CMTEB
Method Classification Pairwise Class. Reranking Retrieval Average
Text-Only (Piccolo2) 74.59 90.24 70.00 74.37 70.95
Text-Only (Conan) 75.02 91.64 72.77 76.68 72.64
ASR + Text (Piccolo2) 72.08 86.09 64.43 54.69 60.78
ASR + Text (Conan) 72.67 88.04 66.65 55.13 61.83
Ours (Piccolo2) 70.98 90.18 68.08 65.99 65.95
4 Experiments
4.1 Implementation Details
For text embedding, we utilize piccolo-large-zh-v2 (Huang et al., 2024) as our text encoder, which provides 1024-dimensional text representations. For speech encoding, we employ Whisper-large-v3 encoder (Radford et al., 2023), leveraging its robust speech feature extraction capabilities trained on large-scale speech-text data.
Training Infrastructure and Schedule
The training process is conducted on a large-scale distributed system consisting of 256 NVIDIA V100 32GB GPUs. We use DeepSpeed for distributed training with mixed-precision (FP16) to optimize memory usage and training efficiency. The entire training process spans 168 hours, divided into two stages:
  • Pre-training: 3 epochs with learning rate $1e^{-5}$ and batch size 8 per GPU
  • Fine-tuning: 3 epochs with learning rate $8e^{-6}$ and batch size 8 per GPU
We employ the AdamW (Loshchilov and Hutter, ) optimizer with weight decay of 0.01 and a linear learning rate scheduler with 10% warmup steps. The temperature parameter $\tau$ in contrastive learning is set to 0.07.
4.2 Datasets
In the training process, different datasets are used at each stage. For the first stage, which is the alignment phase, we collected publicly available data (Emilia) (He et al., 2024) and a subset of data from the internet, totaling 170k hours. In the second stage, we used text training data from PiccoloV2, which includes both open-source data and data generated using specific methods. For each query, we generated data using CosyVoice-300M (Du et al., 2024). During the generation process, to ensure the model’s generalization ability, we also used six different voice tones for generation. In the second stage, a total of 80k hours of data were used for contrastive learning training. Inspired by Emilia, we also built a data filtering pipeline. In the first stage of data filtering, we applied several techniques, including Source Separation, Speaker Diarization, and Fine-grained Segmentation by VAD, to split the speech into clean speech from a specific speaker. Finally, we used ASR to obtain the corresponding text. For both the first and second stages of data, we employed DNSMOS P.835 OVRL(Reddy et al., 2022), keeping only the speech with scores above 3.0. Additionally, any speech segments with average phoneme durations exceeding 1.5 times the interquartile range (IQR) were discarded.
4.3 Results
4.3.1 CMTEB
As an authoritative benchmark for evaluating text embedding tasks, MTEB (Massive Text Embedding Benchmark) (Muennighoff et al., 2022) was introduced by Muennighoff et al. in 2022. For Chinese scenarios, Xiao et al. developed CMTEB (Chinese Massive Text Embedding Benchmark) (Xiao et al., 2024) in 2023, which contains 35 datasets covering 6 major categories: Classification, Clustering, Pair Classification, Rerank, Retrieval, and Semantic Textual Similarity (STS). In our experiments, we utilized CosyVoice text-to-speech technology to convert these textual data into corresponding speech data for evaluating our model’s performance. We show the results in Table 1. As shown in the table, our method outperforms the two-stage ASR approach, regardless of whether Piccolo or the current state-of-the-art method, Conan (Li et al., 2024), is used for alignment. We believe that if the alignment text model uses Conan, the performance is likely to improve further.
4.3.2 More Dataset
To evaluate our model, we conducted Top-1 and Top-3 accuracy tests on a knowledge base containing tens of thousands of entries across multiple domains. From the perspectives of both time efficiency and accuracy, the speech embedding model from the first stage demonstrated significant advantages. Overall, the accuracy of our approach falls between the ASR pipeline and text-only models. Compared to the two-stage method, our approach improves speed by more than 50% and accuracy by around 8%. The results are shown in Table 2.
Table 2: Evals of Different Methods
Method Time(s) Top1-Acc. Top3-Acc.
Text-Only 0.03 90.41 95.89
ASR Pipeline 0.67 79.45 85.62
Project to text 0.43 24.49 54.08
Align with CTC 0.31 78.08 84.25
Ours 0.31 86.36 92.47
Analysis
During our investigation, we explored several alternative approaches that provided valuable insights despite their limitations.
The first approach (Project to text) follows a design philosophy similar to LLAVA models, projecting speech features into text space through an adapter before feeding them into BERT. While this method yields some results, it performs the worst among all baseline approaches. In the process, we even experimented with larger-parameter MLP layers and multi-stage freezing training strategies, but none achieved satisfactory outcomes. Although it leverages pre-trained speech models, it fails to effectively align the speech and text embedding models. We also explored alignment strategies similar to those used in Qwen-Audio, but the results were similarly suboptimal. We believe that the significant difference in sequence length between speech and text poses a fundamental challenge in embedding training tasks, making it difficult to directly integrate with text embedding models through cascaded transformations. Finding an effective way to better align the sequence lengths of speech and text may be the key to overcoming the limitations of this approach.
We also explored using Connectionist Temporal Classification (CTC) loss for temporal alignment between speech and text sequences. While CTC provides explicit temporal supervision and theoretically facilitates better alignment between acoustic patterns and semantic content, its strong focus on frame-level alignment proves to be less effective for semantic retrieval tasks. The rigid frame-by-frame correspondence enforced by CTC restricts the model’s ability to learn flexible, context-aware representations that are critical for robust cross-modal retrieval. Our experiments reveal that this approach performs particularly poorly on longer speech segments, where maintaining long-range dependencies becomes essential.
4.4 Ablation Study
For experimental convenience, we conducted ablation experiments on training methods and speech pretraining models using a sampled dataset with 10,000 hours of data.
Table 3: Ablation Study
Stage Pretrain Top1-Acc. Top3-Acc.
Only 1 Whisper 50.68 61.67
Only 2 Hubert 13.70 21.23
SenseVoice-small 10.96 20.55
Whisper 32.19 42.17
All Whisper 59.59 72.60
From the experimental results in Table 3, it can be observed that whether only Stage 1 or Stage 2 is used, the performance of the final model decreases significantly. Moreover, directly performing contrastive training on a model without Stage 1 alignment poses significant challenges. We believe this is because a speech pretraining model tends to focus more on acoustic information. Without aligning the speech features with textual information in the first stage, it becomes challenging to extract meaningful speech representations solely through contrastive learning. Therefore, the alignment process in the first stage is crucial.
Additionally, in the experiments focusing solely on Stage 2 training, we compared three pretraining models: Hubert (Hsu et al., 2021), SenseVoice-small (FunAudioLLM, 2024), and Whisper (Radford et al., 2023). The results showed that Whisper outperformed the other two significantly, likely due to its use of a large amount of speech data during pretraining.
5 Conclusion
In this paper, we present a novel end-to-end speech-text embedding model designed to address the challenges of high latency and error propagation that are common in traditional sequential architectures. Our approach not only enhances the integration of RAG techniques within SLLMs, but also marke a significant advancement in real-time speech interaction systems. We provide a comprehensive overview of the methods employed in our model, highlighting their effectiveness in optimizing the alignment of speech and text modalities. Moreover, we observe a growing trend in utilizing discrete tokens as end-to-end inputs for SLLMs, which we believe presents a promising avenue for future research in the development of more efficient and capable speech-text embedding models.